# Coercivity Panorama of Dynamic Hysteresis

### v1

@@ Abstract
We study the stochastic $\phi^{4}$ model under periodic driving by an external field $H$ at different scales of driving rate $v_{H}$, where the noise strength $\sigma$ quantifies the deviation of the system size from the thermodynamic limit. For large systems with small $\sigma$, we find the coercivity $H_{c}=H(\langle\phi\rangle=0)$ sequentially exhibits distinct behaviors with increasing $v_{H}$: $v_{H}$-scaling increase from zero, stable plateau ($v_{H}^{0}$), $v_{H}^{1/2}$-scaling increase, and abrupt decline to disappearance. The $H_{c}$-plateau reflects the competition between thermodynamic and quasi-static limits, namely, $\lim_{\sigma\to 0}\lim_{v_{H}\to 0}H_{c}=0$, and $\lim_{v_{H}\to 0}\lim_{\sigma\to 0}H_{c}=H^{*}$. Here, $H^{*}$ is exactly the first-order phase transition (FOPT) point. In the post-plateau slow-driving regime, $H_{c}-H^{*}$ scales with $v_{H}^{2/3}$. Moreover, we reveal a finite-size scaling for the coercivity plateau $H_{P}$ as $(H^{*}-H_{P})\sim\sigma^{4/3}$ by utilizing renormalization-group theory. These predicted scaling relations are demonstrated in magnetic hysteresis obtained with the Curie-Weiss model. Our work provides a panoramic view of the finite-time evolution of the stochastic $\phi^{4}$ model, bridges dynamics of FOPT and dynamic phase transition, and offers new insights into finite-time/finite-size effect interplay in non-equilibrium thermodynamics.
@@
Figure 1: Hysteresis loops and coercivity panorama of the stochastic $\phi^{4}$ model. Hysteresis loops at (a): $v_{H}=10^{-5}$(solid), $4\times 10^{-4}$(dashed); (b): $v_{H}=0.01$(solid), $0.02$(dashed), $0.04$(dash-dotted); (c): $v_{H}=10$(solid), $20$(dashed), $40$(dash-dotted); (d): $v_{H}=100$(solid), $500$(dashed), $1500$(dash-dotted). (e) Coercivity $H_{c}$ as a function of driving rate $v_{H}$ with $\sigma=0.7$. (f) Coercivity landscape $H(v_{H},\sigma)$, where the noise strength from top to bottom is $\sigma=$ 0, 0.1, 0.2, 0.3, 0.4, 0.5, 0.6, 0.7, 0.8, 0.9, 1.0, 2.0. (g) A contour illustration of (f) after interpolation. In this plot, limited driving amplitude $H_{m}=50$ is applied in (a-e) while (f-g) not limited driving amplitude to ensure that the order parameter is always saturated; $a_{2}=-4$, and all quantities are nondimensionalized by $a_{4}$ and $\lambda$.
@@
Introduction.—Hysteresis is ubiquitous across natural and engineered systems, spanning micro- to macro-scales. This phenomenon manifests in diverse fields, including magnetic and mechanical materials [1, 2, 3, 4, 5], electric and optical devices [6, 7], thermal systems [8, 9, 10], biological cycles [11], chemical reactions [12], and even economies [13]. Characterized by history-dependent dynamics, hysteresis reveals a system’s delayed response to external perturbations, where its current state relies on both instantaneous driving parameters and their temporal evolution. Within the realm of physics, magnetic hysteresis has garnered country-long attention due to its rich phenomenology, with extensive experimental and theoretical investigations [1, 14, 15, 16, 17]. It reveals fundamental magnetic properties—anisotropy, coercivity, and remanence—that underpin advancements in magnetic technologies, including permanent-magnet devices [18] and magnetic random access memories [19].
@@
Notably, magnetic hysteresis loops exhibit diverse morphologies—some display steep characteristics indicative of first-order phase transitions (FOPT), while others feature smoother, thinner profiles [20, 21, 22]. A central focus in hysteresis research lies in how hysteresis loops evolve with driving parameters, for example, the rate-dependent deformations in Fig. 1(a-d). This issue encompasses not only the scaling laws of the hysteresis loop area with dynamic parameters [23, 24, 25, 26, 27, 28, 29] but also delves into phenomena like the dynamic phase transitions (DPT) [30, 31, 32]. Critical scaling analyses reveal a fundamental dichotomy: systems with and without quasi-static hysteresis exhibit distinct dynamical behaviors [23, 24], rooted in contrasting ergodic properties [33]. This disparity underscores the interplay between the system’s relaxation capability and driving dynamics [34, 35, 36, 37, 38, 39, 40, 41]. While mean-field theories predict a $2/3$ power law scaling for finite-time deviation from the quasi-static hysteresis in slow-driving FOPT systems [23, 42], experiments and finite-dimensional Ising models show a richer diversity of scaling behaviors [43, 44, 45, 46, 47, 48, 49, 50, 51].
@@
Despite these significant advancements, current studies on hysteresis confront three critical challenges: i) In microscopic theories rooted in nonequilibrium statistical physics, scaling relations of dynamic hysteresis across diverse time scales remain largely unexplored [23, 52, 42, 53, 54]; ii) For phenomenological models in magnetism, such as the Jiles-Atherton model [55, 56, 57, 58], Preisach model [59, 60, 61, 62] and Stoner-Wohlfarth model [63, 64], there is a significant lack of understanding of the continuous transition between quasi-static and dynamic hysteresis [55, 65]. This raises a critical question: do hysteresis loops have a well-defined quasi-static limit, and how does it emerge with decreasing driving rate? iii) The comparison between experimentally measured and theoretically predicted scaling laws for hysteresis suffers from ambiguity [48, 66, 9, 10], primarily due to incomplete characterization of hysteresis dependence on driving rate across the entire accessible frequency regimes.
@@
To address these gaps, we introduce the concept of coercivity panorama[Fig. 1(e)], derived from the driven stochastic $\phi^{4}$ model. This framework offers a unified lens to reconcile scaling relations across time scales and system sizes, highlighting competition between thermodynamic and quasi-static limits in interacting systems. Concomitantly, the derived order parameter’s evolution equation serves as a natural bridge between microscopic dynamics and macroscopic hysteretic behaviors, resolving long-standing theory-experiment mismatches in hysteresis modeling.
@@
where the constants $a_{2}$ and $a_{4}$ encode the reduced temperature and intrinsic interaction strength, respectively. For $a_{2}<0$ at zero field ($H=0$), the free energy landscape exhibits $\mathbb{Z}_{2}$-symmetric bistability. Under weak applied fields satisfying $|H|<|H_{\mathrm{sp}}^{\pm}|$, this double-well potential becomes tilted while maintaining both local minima. Here, $H_{\mathrm{sp}}^{\pm}\equiv\pm H^{*}\equiv\pm\sqrt{(4/27)(-a_{2}^{3}/a_{4})}$ denotes the spinodal field, at which the metastable state ceases to exist and the metastable minimum merges with the unstable maximum at $\phi_{\mathrm{sp}}^{\pm}=\pm\phi^{*}\mp\text{sgn}(H_{\mathrm{sp}}^{\pm})\sqrt{-a_{2}/(3a_{4})}$. The superscripts $\pm$ denote transitions toward positive and negative $\phi$, respectively. The dynamics of $\phi=\phi(t)$ under a time-dependent field $H=H(t)$ can be described by the Langevin equation [68, 48]
@@
where $\lambda$ sets a reference timescale and $\zeta(t)$ is a Gaussian white noise with amplitude $\sqrt{2\lambda\sigma^{2}}$ satisfying $\langle\zeta(t)\rangle=0$ and $\langle\zeta(t)\zeta(t^{\prime})\rangle=2\lambda\sigma^{2}\delta(t-t^{\prime})$, and $\sigma\geq 0$ denoting the noise strength. The system defined by Eqs. (1) and (2) is referred to as stochastic $\phi^{4}$ model throughout this work. In the deterministic limit ($\sigma\to 0$), the model describes a sharp FOPT [23, 53], exactly occurring at the spinodal fields $H=H_{\mathrm{sp}}^{\pm}$ under quasi-static driving between saturation fields of opposite signs. Finite driving rates delay the transition beyond $H=H_{\mathrm{sp}}^{\pm}$ due to dynamic hysteresis [23], while finite fluctuations ($\sigma>0$), representing deviations from the thermodynamic limit in our framework, cause premature escape from metastable states before reaching $H=H_{\mathrm{sp}}^{\pm}$. The competition between these finite-time and finite-size effects leads to complex nonequilibrium dynamics.
To directly capture the dynamics of the stochastic $\phi^{4}$ model, we first simulate its evolution with a discretized form of Eq. (2) (See Supplemental Materials [69] for details). Figures 1(a-d) display hysteresis loops for four different driving regimes, generated under a field swept triangularly between ±$\pm H_{m}$ at a rate $dH/dt=\lambda v_{H}$, where $H_{m}>0$ is the amplitude. Two key features emerge from these figures: (i) The hysteresis loop is symmetric with respect to the origin and the loop width increases monotonically with $v_{H}$ when the saturation field is smaller than $H_{m}$; however, at very fast driving rates, the loops shrink and form narrow cycles centered around the initial value of the order parameter. (ii) In the first two regimes, the hysteresis loops exhibit clear signatures of FOPT, characterized by a distinct turning point that triggers a rapid phase transition. In contrast, this feature becomes weaker and even vanishes in the latter two regimes.
@@
which is clearly illustrated by the contour plot in Fig. 1(g). Taking the quasi-static limit first implies that finite fluctuations are always present to ensure ergodicity. In contrast, taking the thermodynamic limit first eliminates fluctuations, causing the system to remain trapped in metastable states even under quasi-static driving. This distinction has also been reflected in the difference between deterministic mean-field results and fluctuating Monte Carlo simulations of the Ising model [45].
Finite-time scaling relations.—For the stochastic $\phi^{4}$ model under a linearly driven external field, the corresponding Fokker-Planck equation of Eq. (2) reads
@@
where $P(\phi,t)$ denotes the probability density of the system having the value $\phi$ at time $t$. Multiplying Eq. (4) by $\phi$ and integrating the resulting expression over all possible values of $\phi$, we obtain an ensemble-averaged equation,
@@
where $dH=\lambda v_{H}dt$ has been used. This equation serves as the first main result of this Letter, capturing the evolution of measurable macroscopic quantities and linking microscopic stochastic models with phenomenological equations widely used in experimental magnetism [15].
@@
Here, the first term on the right-hand side is related to the slope of the $\langle\phi\rangle-H$ curve at the coercivity point; the second term characterizes the skewness of the probability distribution of $\phi$ when $\langle\phi\rangle=0$. As a result of Eq. (6), three distinct scaling regimes of coercivity with respect to the driving rate are identified(See [69] for detailed derivations), as illustrated in Tab. 1.
Table 1: Finite-time coercivity scaling relations for the stochastic $\phi^{4}$ model.
near-equilibrium post-plateau fast-driving
$H_{c}\sim v_{H}$ $H_{c}-H_{P}\sim(v_{H}-v_{P})^{2/3}$ $H_{c}\sim v_{H}^{1/2}$
In the near-equilibrium regime, for non-divergent equilibrium value of $\evaluated{d\langle\phi\rangle/dH}_{\langle\phi\rangle=0}$, the scaling $H_{c}\sim v_{H}$ is denoted by the dashed line in Fig. 1(f). This linear response relation is universal for systems weakly driven out of equilibrium [71, 72, 73]. In the fast-driving regime, the dynamics is effectively governed by a passive relaxation process toward the mono-stable minimum induced by the continuously increasing field. The interplay between the relaxation timescale and the field variation rate leads to a $H_{c}\sim v_{H}^{1/2}$ scaling, which is denoted by the dash-dotted line in Fig. 1(f). While previous analytical studies on mean-field models often overlooked this fast-driving regime, this scaling relation finds support in existing numerical [74] and experimental evidence [75].
@@
where $H_{P}$ denotes the height of the coercivity plateau, $H^{*}$ is that height in the deterministic limit, and $v_{P}$ is the reference driving rate at which the coercivity plateau is observed [77]. Up to now, the noise strength has been treated as a phenomenological parameter reflecting finite-size effects. The explicit dependence of $\sigma$ on system size $N$ can be illustrated through specific microscopic examples [47, 54].
Figure 2: Reduced coercivity after the FOPT plateau. (a) Coercivity reduced by the reference plateau coercivity value. Curves from right to left correspond to noise strengths $\sigma=$ 1.0, 0.8, 0.7, 0.5, 0.4, 0.3. (b) Coercivity reduced by the minimum observed coercivity value within the simulation timescale shown in Fig. 1(f). Curves from left to right correspond to $\sigma=$ 0.3, 0.2, 0.1, 0. Parameters: $a_{2}=-4$. All quantities are nondimensionalized by $a_{4}$ and $\lambda$.
Finite-size corrections to the $2/3$ power law scaling.—Unlike the universal scaling observed in the near-equilibrium and fast-driving regimes, the scaling behavior near the coercivity plateau is highly sensitive to system size and constrained by the finite timescales accessible in both simulations and experiments. By expanding the dynamic equation (2) around the coercivity plateau point $(v_{P},H_{P})$, we obtain a finite-time scaling as $H_{c}-H_{P}\sim(v_{H}-v_{P})^{2/3}$ [69], where both $v_{P}$ and $H_{P}$ depend on system sizes (8). In the deterministic limit $\sigma\to 0$, $H_{P}\rightarrow H^{*}$ and $v_{P}\rightarrow 0$, such that this scaling reduces to $H_{c}-H^{*}\sim v_{H}^{2/3}$. This is consistent with the scaling, $A-A_{0}\sim v_{H}^{2/3}$, characterized by subtracting the quasistatic hysteresis loop area $A_{0}$ from the finite-time area $A$ [23, 42, 54], because the loop geometry directly leads to $A-A_{0}\propto H_{c}-H^{*}$. This $2/3$-exponent scaling for finite $\sigma$ is demonstrated by the dotted lines in Fig. 2(a). However, for very small noise strengths, the genuine plateau point becomes unreachable within accessible simulation timescales [as shown in the upper set of Fig. 1(f)], making it challenging to clearly characterize the coercivity plateau. In such cases, it is natural to take the minimum observed coercivity $(H_{c}^{\mathrm{min}})$ as a quasi-static value [48]. Fig. 2(b) illustrates the scaling relation $H_{c}-H_{c}^{\mathrm{min}}$ with respect to $v_{H}$, which indicates a transient $2/3$-exponent scaling. This localized scaling behavior has also been reported in Ref. [54]. Depending on the accessible rate range and specific models, the prominent effective scaling exponent can be either smaller or larger than $2/3$ [48, 78]. This variability could explain observed deviations from deterministic mean-field predictions in the accessible slow-driving regime. In this sense, the global scaling landscape of coercivity provides a more robust and reliable characterization than the localized scaling behavior of the reduced coercivity near the FOPT plateau.
@@
Magnetic coercivity in Curie-Weiss model.—To illustrate the finite-size scaling obtained from the stochastic $\phi^{4}$ model, we focus on the magnetic hysteresis of the Curie-Weiss model [79, 80, 54, 81], which is a solvable model of magnetism, also known as the mean-field Ising model (See End Matters for details). By comparing the roles of $\sigma$ and $N$ in the Fokker-Planck equations of the stochastic $\phi^{4}$ and Curie-Weiss models, one has $\sigma\sim N^{-1/2}$. By rescaling the axes in the $H_{c}-v_{H}$ plot according to the scaling relations in Eq. (8), it is found that the coercivity curves for different system sizes collapse onto a single universal curve, as shown in Fig. 3(a), confirming the validity of the finite-size scaling relations. This universal curve diverges at a critical driving rate that yields $H_{c}=H^{*}$. Here $H^{*}$ is the spinodal field of the Curie-Weiss model. In the infinite-size limit, this critical driving rate approaches zero, rendering the region to its left inaccessible. This demonstrates that the coercivity plateau resulting from taking the thermodynamic limit ($\sigma\rightarrow 0$) first in Eq.(3) does not vanish in the slow-driving regime. In contrast, for finite-size systems, coercivity values span both sides of this critical point, with the gently varying region at lower driving rates reflecting the presence of a coercivity plateau.
On the other hand, it is worth mentioning that the $v_{H}$-scaling in the near-equilibrium regime and the finite-size scaling relations associated with the FOPT plateau are universal across both the stochastic $\phi^{4}$ and Curie-Weiss models. However, in the fast-driving regime, despite sharing similar physical mechanisms, the specific scaling differs between the two models due to fundamental differences in their full dynamical equations. In the Curie-Weiss model, as shown in Fig. 3(b), the scaling behavior (yellow dash-dotted curve) in the fast-driving regime follows the Lambert W-function $W(v_{H})$ [69]. Moreover, between the FOPT plateau and the fast-driving regime, an intermediate scaling (dashed line), $H_{c}\sim v_{H}^{1/3}$ [69], emerges. While this scaling is absent as a sustained regime in the stochastic $\phi^{4}$ model, it is briefly passed through without being maintained over an observable timescale. This regime bridges the region governed by FOPT characteristics and the passive relaxation dynamics at high driving rates. These discrepancies highlight the sensitivity of far-from-equilibrium dynamics to model-specific details, even when the underlying physical picture remains similar. Consequently, different microscopic models may reveal more varied scaling relations in this regime  [52].
Figure 3: Coercivity of the Curie-Weiss model as a function of driving rate. (a) Finite-size scaling collapse of coercivity plateau. (b) Detailed scalings in the relatively fast-driving regimes after the plateau, where $W(v_{H})$ denotes the Lambert W function (real-valued branch). A bounded driving amplitude $H_{m}=1$ is imposed in (a), while the driving amplitude is unconstrained to ensure full saturation of the order parameter in (b). In this plot, we choose $\tau_{0}=2,\beta=2,J=1$.
@@
Conclusion and discussion.—We have introduced the concept of the coercivity panorama to characterize the irreversibility of hysteresis phenomena in interacting systems across different time scales. In the periodically driven stochastic $\phi^{4}$ model, we analytically obtained the finite-time and finite-size scaling of the coercivity. The coercivity panorama bridges the gap between irreversibility in the linear response regime and irreversibility with signatures of FOPT, reflecting the competition between the thermodynamic limit and the quasi-static limit. Remarkably, the scaling relations near the coercivity plateau are universal, which is confirmed in the Curie-Weiss model. However, in the fast-driving regime, where the system transitions from the finite-time behavior of FOPT to the DPT regime, the stochastic $\phi^{4}$ and Curie-Weiss models exhibit different scaling behaviors.

### v2

@@ Abstract
We study the stochastic $\phi^{4}$ model under periodic driving by an external field $H$ at different scales of driving rate $v_{H}$, where the noise strength $\sigma$ quantifies the deviation of the system size from the thermodynamic limit. For large systems with small $\sigma$, we find the coercivity $H_{c}\equiv H(\langle\phi\rangle=0)$ sequentially exhibits distinct behaviors with increasing $v_{H}$: $v_{H}$-scaling increase from zero, stable plateau ($v_{H}^{0}$), $v_{H}^{1/2}$-scaling increase, and abrupt decline to disappearance. The $H_{c}$-plateau reflects the competition between thermodynamic and quasi-static limits, namely, $\lim_{\sigma\to 0}\lim_{v_{H}\to 0}H_{c}=0$, and $\lim_{v_{H}\to 0}\lim_{\sigma\to 0}H_{c}=H^{*}$. Here, $H^{*}$ is exactly the first-order phase transition (FOPT) point. In the post-plateau slow-driving regime, $H_{c}-H^{*}$ scales with $v_{H}^{2/3}$. Moreover, we reveal a finite-size scaling for the coercivity plateau $H_{P}$ as $(H^{*}-H_{P})\sim\sigma^{4/3}$ by utilizing renormalization-group theory. These predicted scaling relations are demonstrated in magnetic hysteresis described by the $N$-spin Curie-Weiss model where $N^{-1/2}\sim\sigma$. Our work provides a panoramic view of finite-time scalings of the hysteresis phenomenon, bridges dynamics of FOPT and dynamic phase transition, and offers new insights into finite-time/finite-size effect interplay in non-equilibrium thermodynamics.
@@
Figure 1: Hysteresis loops and coercivity panorama of the stochastic $\phi^{4}$ model. Hysteresis loops at (a): $v_{H}=10^{-5}$(solid), $4\times 10^{-4}$(dashed); (b): $v_{H}=0.01$(solid), $0.02$(dashed), $0.04$(dash-dotted); (c): $v_{H}=10$(solid), $20$(dashed), $40$(dash-dotted); (d): $v_{H}=100$(solid), $500$(dashed), $1500$(dash-dotted). (e) Coercivity $H_{c}$ as a function of driving rate $v_{H}$ with $\sigma=0.7$. (f) Coercivity landscape $H(v_{H},\sigma)$, where the noise strength from top to bottom is $\sigma=$ 0, 0.1, 0.2, 0.3, 0.4, 0.5, 0.6, 0.7, 0.8, 0.9, 1.0, 2.0. (g) A contour illustration of (f) after interpolation. In this plot, limited driving amplitude $H_{m}=50$ is applied in (a-e) while (f-g) not limited driving amplitude to ensure that the order parameter is always saturated; $a_{2}=-4$, and all quantities are nondimensionalized by $a_{4}$ and $\lambda$. FOPT, first-order phase transition; DPT, dynamic phase transition.
@@
Introduction.— Hysteresis is ubiquitous across natural and engineered systems, spanning micro- to macro-scales. This phenomenon manifests in diverse fields, including magnetic and mechanical materials [1, 2, 3], electric and optical devices [4, 5, 6], thermal systems [7, 8, 9], biological cycles [10], chemical reactions [11], and even economies [12]. Characterized by history-dependent dynamics, hysteresis reveals a system’s delayed response to external perturbations, where its current state relies on both instantaneous driving parameters and their temporal evolution. Within the realm of physics, in particular, magnetic hysteresis has garnered country-long attention due to its rich phenomenology, with extensive experimental and theoretical investigations [13, 14, 15, 16, 17, 18]. It reveals fundamental properties like anisotropy, coercivity, and remanence, which underpin advancements in magnetic technologies, including permanent-magnet devices [19] and magnetic random access memories [20].
@@
Notably, hysteresis loops exhibit diverse morphologies—some display steep characteristics indicative of first-order phase transitions (FOPT), while others feature smoother, thinner profiles [21, 22, 23]. A central focus in hysteresis research lies in how hysteresis loops evolve with driving parameters, for example, the rate-dependent deformations in Fig. 1(a-d). This issue encompasses not only the scaling laws of the hysteresis loop area with dynamic parameters [24, 25, 26, 27, 28, 29, 30] but also delves into phenomena like the dynamic phase transitions (DPT) [31, 26, 32, 33]. Critical scaling analyses reveal a fundamental dichotomy: systems with and without quasi-static hysteresis exhibit distinct dynamical behaviors [24, 25], rooted in contrasting ergodic properties [34]. This disparity underscores the interplay between the system’s relaxation capability and driving dynamics [35, 36, 37]. While mean-field theories predict a $2/3$ power law scaling for finite-time deviation from the quasi-static hysteresis in slow-driving FOPT systems [24, 38], experiments and finite-dimensional Ising models show a richer diversity of scaling behaviors [39, 40, 41, 42, 43, 44, 45, 46, 47].
@@
Despite these significant advancements, current studies on hysteresis confront three critical challenges: i) In microscopic theories rooted in nonequilibrium statistical physics, scaling relations of dynamic hysteresis across diverse time scales remain largely unexplored [24, 48, 38, 49, 50]; ii) For phenomenological models in magnetism, such as the Jiles-Atherton model [51, 52], Preisach model  [53, 54] and Stoner-Wohlfarth model [55], there is a significant lack of understanding of the continuous transition between quasi-static and dynamic hysteresis [51, 56]. This raises a critical question: do hysteresis loops have a well-defined quasi-static limit, and how does it emerge with decreasing driving rate? iii) The comparison between experimentally measured and theoretically predicted scaling laws for hysteresis suffers from ambiguity [9, 8, 57, 46], primarily due to incomplete characterization of hysteresis dependence on driving rate across the entire accessible frequency regimes.
@@
To address these gaps, we introduce the concept of coercivity panorama [Fig. 1(e)], derived from the driven stochastic $\phi^{4}$ model. This framework offers a unified lens to reconcile scaling relations across time scales and system sizes, highlighting competition between thermodynamic and quasi-static limits in interacting systems. Concomitantly, the derived order parameter’s evolution equation serves as a natural bridge between microscopic dynamics and macroscopic hysteretic behaviors, providing a framework for reconciling long-standing theory-experiment mismatches in hysteresis modeling.
@@
where the constants $a_{2}$ and $a_{4}$ encode the reduced temperature and intrinsic interaction strength, respectively. For $a_{2}<0$ at zero field ($H=0$), the free energy landscape exhibits $\mathbb{Z}_{2}$-symmetric bistability. Under weak applied fields satisfying $|H|<|H_{\mathrm{sp}}^{\pm}|$, this double-well potential becomes tilted while maintaining both local minima. Here, $H_{\mathrm{sp}}^{\pm}\equiv\pm H^{*}\equiv\pm\sqrt{(4/27)(-a_{2}^{3}/a_{4})}$ denotes the spinodal field, at which the metastable state ceases to exist and the metastable minimum merges with the unstable maximum at $\phi_{\mathrm{sp}}^{\pm}\equiv\pm\phi^{*}\equiv\mp\sqrt{-a_{2}/(3a_{4})}$. The superscripts $\pm$ denote transitions toward positive and negative $\phi$, respectively. The dynamics of $\phi=\phi(t)$ under a time-dependent field $H=H(t)$ can be described by the Langevin equation [59, 46]
@@
where $\lambda$ sets a reference timescale and $\zeta(t)$ is a Gaussian white noise with amplitude $\sqrt{2\lambda\sigma^{2}}$ satisfying $\langle\zeta(t)\rangle=0$ and $\langle\zeta(t)\zeta(t^{\prime})\rangle=2\lambda\sigma^{2}\delta(t-t^{\prime})$, and $\sigma\geq 0$ denoting the noise strength. The system defined by Eqs. (1) and (2) is referred to as stochastic $\phi^{4}$ model throughout this work. In the deterministic limit ($\sigma\to 0$), the model describes a sharp FOPT [24, 49], exactly occurring at the spinodal fields $H=H_{\mathrm{sp}}^{\pm}$ under quasi-static driving between saturation fields of opposite signs. Finite driving rates delay the transition beyond $H=H_{\mathrm{sp}}^{\pm}$ due to dynamic hysteresis [24], while finite fluctuations ($\sigma>0$), representing deviations from the thermodynamic limit in our framework [60, 61], cause premature escape from metastable states before reaching $H=H_{\mathrm{sp}}^{\pm}$. The competition between these finite-time and finite-size effects leads to complex nonequilibrium dynamics.
To directly capture the dynamics of the stochastic $\phi^{4}$ model, we first simulate its evolution with a discretized form of Eq. (2) (See companion paper [62] for details). Figures 1(a-d) display stable hysteresis loops in four different driving regimes, generated under a field swept triangularly with an amplitude $H_{m}$ and at a rate $|dH/dt|=\lambda v_{H}$. Two key features emerge from these figures: i) The hysteresis loop is symmetric with respect to the origin and the loop width increases monotonically with $v_{H}$ when the saturation field is smaller than $H_{m}$; however, at very fast driving rates, the loops shrink and form narrow cycles centered near the initial value of the order parameter; ii) In the first two regimes, the hysteresis loops exhibit clear signatures of FOPT, characterized by a distinct turning point that triggers a rapid phase transition. In contrast, this feature becomes weaker and even vanishes in the latter two regimes.
@@
which is clearly illustrated by the contour plot in Fig. 1(g). Taking the quasi-static limit first implies that finite fluctuations are always present to ensure ergodicity. In contrast, taking the thermodynamic limit first eliminates fluctuations, causing the system to remain trapped in metastable states even under quasi-static driving, rendering the global equilibrium state unreachable. This distinction has also been reflected in the difference between deterministic mean-field results and fluctuating Monte Carlo simulations of the Ising model [42]. In this sense, for macroscopic and mesoscopic materials, within the achievable observation time, the quasi-static hysteresis loop can be well-defined with the coercivity associated with the FOPT point; for microscopic materials with non-negligible fluctuations, it is expected that the hysteresis loop will disappear during slow field sweeping and approach the equilibrium magnetization curve.
Finite-time scaling relations.—To quantify the evolution of the ensemble-averaged order parameter, we derive the corresponding Fokker-Planck equation of Eq. (2) as
@@
where $P(\phi,t)$ denotes the probability density of the system having the value $\phi$ at time $t$. Multiplying Eq. (4) by $\phi$ and integrating the resulting expression over all possible values of $\phi$, an ensemble-averaged equation is obtained as,
@@
where $dH=\lambda v_{H}dt$ has been used. Hereafter, the increasing branch of the field function is considered in analytical calculations, unless specifically stated otherwise. As the first main result of this Letter, Eq. (5) captures the evolution of measurable macroscopic quantities and connects microscopic stochastic models with phenomenological equations widely used in magnetism [18].
@@
Here, the first term on the right-hand side is related to the slope of the $\langle\phi\rangle-H$ curve at the coercivity point; the second term characterizes the skewness of the probability distribution of $\phi$ when $\langle\phi\rangle=0$. As a result of Eq. (6), three distinct scaling regimes for coercivity with respect to the driving rate are identified: the near-equilibrium regime, the post-plateau regime and the fast-driving regime [62]. In the near-equilibrium regime, for non-divergent equilibrium value of $\evaluated{d\langle\phi\rangle/dH}_{\langle\phi\rangle=0}$, the scaling
$H_{c}\sim v_{H}$ (7)
is denoted by the dashed line in Fig. 1(f). This linear response relation is universal for systems weakly driven out of equilibrium [65, 66, 67, 68, 69]. In the fast-driving regime, the dynamics is effectively governed by a passive relaxation process toward the mono-stable minimum induced by the continuously increasing field. The interplay between the relaxation timescale and the field variation rate leads to a scaling as
$H_{c}\sim v_{H}^{1/2},$ (8)
which is denoted by the dash-dotted line in Fig. 1(f). While previous analytical studies on mean-field models often overlooked this fast-driving regime, this scaling finds support in existing numerical [70] and experimental evidence [71].
@@
where $H_{P}$ denotes the height of the coercivity plateau, $H^{*}$ is that height in the deterministic limit, and $v_{P}$ is the reference driving rate at which the coercivity plateau is observed [73]. By rescaling the axes in Fig. 1(f) according to the scaling relations in Eq. (10), it is found that the coercivity curves for small $\sigma\in[0.1,1]$ collapse onto a single universal curve, as shown in Fig. 2(a), confirming the validity of the finite-size scaling relations. This universal curve diverges at a critical driving rate that yields $H_{c}=H^{*}$. In the $\sigma\to 0$ limit, this critical driving rate approaches zero, rendering the region to its left inaccessible. This demonstrates that the coercivity plateau resulting from taking the thermodynamic limit first in Eq. (3) does not vanish in the slow-driving regime. In contrast, for finite-size systems, coercivity values span both sides of this critical point, with the gently varying region at lower driving rates reflecting the presence of a coercivity plateau. Up to now, the noise strength has been treated as a phenomenological parameter reflecting finite-size effects. The explicit dependence of $\sigma$ on system size can be determined through specific microscopic examples [45, 50], as we demonstrate in the companion paper [62]
Figure 2: (a) Finite-size scaling of the FOPT plateau. $\sigma$ from 0.1 to 1.0 with interval 0.1. (b) Coercivity reduced by the reference plateau coercivity value. Curves from right to left correspond to noise strengths $\sigma=$ 1.0, 0.8, 0.7, 0.5, 0.4, 0.3. (c) Coercivity reduced by the minimum observed coercivity value within the simulation timescale shown in Fig. 1(f). Curves from left to right correspond to $\sigma=$ 0.3, 0.2, 0.1, 0.
Finite-size corrections to the $2/3$ power law scaling.—Adopting a procedure analogous to the finite-time scaling analysis in the thermodynamic limit [24, 49], we expand the evolution equation (4) around the coercivity plateau point $(v_{P},H_{P})$ to reveal the following finite-time scaling relation [62]
$H_{c}-H_{P}\sim(v_{H}-v_{P})^{2/3}$ (11)
where both $v_{P}$ and $H_{P}$ depend on system sizes [Eq. (10)]. In the deterministic limit ($\sigma\to 0$), $(H_{P},v_{P})\rightarrow(H^{*},0)$, such that this scaling reduces to $H_{c}-H^{*}\sim v_{H}^{2/3}$. This is consistent with the scaling $A-A_{0}\sim v_{H}^{2/3}$, expressed by subtracting the quasistatic hysteresis loop area $A_{0}$ from the finite-time area $A$ [24, 38, 50]. Such a consistency is due to the fact that the hysteresis loop in this regime [Fig. 1(b)] is geometrically approximated to a parallelogram with $A-A_{0}\propto H_{c}-H^{*}$. The $2/3$-exponent scaling for finite $\sigma$ is demonstrated by the dotted lines in Fig. 2(b). However, for very small noise strengths, the genuine plateau point becomes unreachable within accessible simulation timescales [as shown in the upper set of Fig. 1(f)], making it challenging to clearly characterize the coercivity plateau. In such cases, it is practical to take the minimum observed coercivity $(H_{c}^{\mathrm{min}})$ as a quasi-static value [46]. Figure 2(c) illustrates the scaling of $H_{c}-H_{c}^{\mathrm{min}}$ with respect to $v_{H}$, which indicates a transient $2/3$-exponent scaling. This localized scaling behavior has also been reported in Ref. [50]. Depending on the accessible rate range and specific models, the prominent effective scaling exponent can be either smaller or larger than $2/3$ [46, 47]. This variability could explain observed deviations from deterministic mean-field predictions in the accessible slow-driving regime. In this sense, the global scaling landscape of coercivity provides a more robust and reliable characterization than the localized scaling behavior of the reduced coercivity near the FOPT plateau.
@@
Concluding remarks.—We have introduced the concept of the coercivity panorama to characterize the irreversibility of hysteresis phenomena in interacting systems across different timescales. The main findings are summarized in Fig. 3, where the coercivity panorama of the Curie-Weiss model is investigated in the companion paper [62]. The coercivity panorama bridges the gap between irreversibility in the linear response regime and irreversibility with signatures of FOPT, reflecting the competition between the thermodynamic and quasi-static limits. Remarkably, the scaling relations near the coercivity plateau are universal across both the stochastic $\phi^{4}$ and Curie-Weiss models [74]. However, these two models exhibit different scaling behaviors in the fast-driving regime, where the system transitions from the finite-time behavior of FOPT to the DPT regime. The discrepancy in the fast-driving regime highlights the sensitivity of far-from-equilibrium dynamics to model-specific details, even when the underlying physical picture remains similar. Consequently, different microscopic models may reveal more varied scaling relations in this regime  [48, 45].
Figure 3: Finite-time and finite-size scaling relations for saturated hysteresis in interacting systems. The results for the Curie-Weiss model of $N$-spins are obtained in the companion paper [62].